OpenStudy (anonymous):

Which point lies in the solution set for the following system of inequalities? y > x + 4 y > -2x + 2

OpenStudy (anonymous):

A. (-4, 0) B. (-3, 2) C. (2, -1) D. (0, 6)

OpenStudy (ajprincess):

plug in the given points and see which satisfy both the inequalities. let me show one. let us take the first option. (-4,0) y=-4,x=0 x+4=4 -4 is not greater than 4 so it does not satisfy the first inequalty. so it is nt the correct one. similarly can u chck the others? @tiffpink

OpenStudy (anonymous):

y = -3, x = 2 what do I do next?

OpenStudy (ajprincess):

what is x+4=?

OpenStudy (anonymous):

4x ?

OpenStudy (ajprincess):

no. x=2 so x+4=2+4=6 getting it?

OpenStudy (anonymous):

sorry I am not getting it

OpenStudy (ajprincess):

oops sorry.

OpenStudy (ajprincess):

i made a big mistake. I am really sorry. the first value is x nd second is y. so let me take the first one explain. (-4,0) then x=-4 and y=0 x+4 in place of x u can put the number -4 and find the value of x+4. -4+4=0 y=0 y must be greater than x+4. but here both are equal. so the point (-4,0) does not satisfy the first inequality y>x+4. getting it?

OpenStudy (ajprincess):

@tiffpink

OpenStudy (anonymous):

@kelliegirl33 can you please help? :/

OpenStudy (anonymous):

y > x + 4 y > - 2x + 2 Now we are gonna plug in the values and see if the inequality is true. Lets take the first answer choice...(-4,0) <--x = - 4 and y = 0 plug those values into the inequalities... y > x + 4 0 > -4 + 4 0 > 0 (not true, 0 is not GREATER then 0...so A is not correct. Next one...(-3,2) y > x + 4 2 > -3 + 4 2 > 1 (this is true) However, we have to check it with the other inequality... y > - 2x + 2 ...(-3,2) 2 > -2(-3) + 2 2 > 6 + 2 2 > 8 (not true) B. (-3,2) will not work because it has to be true for both inequalities Next one...(2, -1) y > x + 4 -1 > 2 + 4 -1 > 6 (not true) C. will not work Last one...(0,6) y > x + 4 6 > 0 + 4 6 > 4 (correct) check the other inequality... y > -2x + 2.....(0,6) 6 > -2(0) + 2 6 > 0 + 2 6 > 2 (correct) As you can see, (0,6) works in both inequalities and is therefore your answer. If you have any questions about what I just did....ask :)